Mandy Feng employs Jay Johnson at a salary of $35,000 a year. Feng is a subject to employer Social Security taxes at a rate of 6.2% and Medicare taxes at a rate of 1.45% on Johnson's salary. In addition, Feng must pay SUTA tax at a rate of 5.4% and FUTA tax at a rate of 0.6% on the first $7,000 of Johnson's salary. Compute the total cost to Feng of employing Johnson for the year. Round your answer to the nearest cent.
Answer #1
Gross salary 35000.00
Social Security tax 2170.00 =35000*6.2%
Medicare tax 507.50 =35000*1.45%
SUTA tax 378.00 =7000*5.4%
FUTA tax 42.00 =7000*0.6%
Total cost to Feng of employing Johnson 38097.50